OpenStudy (anonymous):

Please help. No one is helping me :( Rewrite the formula. Solve the area of the trapezoid formula for one base. So pretty much make the equation start with b1=

OpenStudy (anonymous):

is there a picture?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

I have to solve for one base, but make the area of the trapezoid equation stay the same.

OpenStudy (anonymous):

well im not quite sure how to figure that out srry

OpenStudy (anonymous):

Well, another similar problem I figured out is solve the slope intercept equation y=mx+b for the slope, m. y=mx+b subtract b to both sides y-b=mx divide by x to both sides y-b/x=m I have to do something like that to the area of the trapezoid formula.

OpenStudy (anonymous):

Area of trape....= 1/2 (b1+b2)h Now u just have to make b1 the subject of the formula, can u do it now?

OpenStudy (anonymous):

i don't know hot to make the b1 alone..

OpenStudy (anonymous):

Rearrange the formula so: 2A = (b1 + b2) h (2A) / h = b1 + b2 b1 = ((2A) / h) - b2